OpenStudy (anonymous):

Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 8^x and y = 2^(x + 2) intersect are the solutions of the equation 8^x = 2^(x + 2). (4 points) Part B Make tables to find the solution to 8^x = 2^(x + 2). Take the integer values of x between -3 and 3. (4 points) Part C: How can you solve the equation 8^x = 2^(x + 2) graphically? (2 points)

OpenStudy (anonymous):

please help i will give a medal out for help

OpenStudy (johnweldon1993):

Well if you were to graph those 2 equations...they would indeed intersect...why? Because at 1 point (1 'x' and 'y' value) they will be equal Lets figure out where... 8^x = 2^(x + 2) 8 can be written as 2^3 right? so we have (2^3)^x = 2^(x + 2) Remember the rules of exponents... 2^(3x) = 2^(x + 2) We just need to figure out when 3x = x + 2 so subtract x from both sides... 2x = 2 x = 1 There is our solution

OpenStudy (anonymous):

what about part b i need help on that part to

OpenStudy (johnweldon1993):

Alright so make a table  All you do is plug in those values for 'x' into your function...and see what they equal